If 33.5 mol of an ideal gas occupies 87.5 L at 13.00 °C, what is the pressure of the gas? pressure:
Answer #1

n 33.5 mo V= Q7-5 L T: 13.00 C T= 286 K P ? Pv=mRT PX 87.5 = 33.5 X 0.0821 X2,86 P &. 989715 atm 6832-2 bas P =